OpenStudy (anna1448):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

U know absolute functions??

OpenStudy (anonymous):

if yes what are the possible values of |x+5|

OpenStudy (anonymous):

English??

OpenStudy (anna1448):

yeah sorry lol i was looking at the question again

OpenStudy (anna1448):

not sure.. im really confused

OpenStudy (anonymous):

|x|=??

OpenStudy (anonymous):

think in terms of >0 and <0

OpenStudy (anna1448):

i really dont get it :( im sorry :(

hartnn (hartnn):

ok, you can use this when \(|x| = y \) then \(\Large x= y \\ \Large or ~ x = -y\)

hartnn (hartnn):

when |x+5| = 4 what can u say ?

OpenStudy (anna1448):

x= -4 ?

hartnn (hartnn):

i mean |x+5| = 4 gives x+5 = 4 OR x+5 =-4

OpenStudy (anna1448):

so then x+5= -4. ?

hartnn (hartnn):

you'll get 2 values of x one from x+5 = 4 and one from x+5 =-4

hartnn (hartnn):

so, x+5 =4 subtract 5 from both sides, what i get ?

hartnn (hartnn):

what you get** ? :P

OpenStudy (anna1448):

1 and 0

OpenStudy (anonymous):

yep

hartnn (hartnn):

you mean -1 ? because 4-5 =-1

OpenStudy (anonymous):

u shud know basic

OpenStudy (anonymous):

how do even get 0 and 1??

OpenStudy (anonymous):

pls show ur calculations

hartnn (hartnn):

x+5-5 = 4-5 0 for 5-5 correct and 1 for 4-5 incorrect

hartnn (hartnn):

4-5 = -1 so you get x = -1 makes sense ?

OpenStudy (anna1448):

listen @cambrige i suck at math i get it. @hartnn yeah that makes more sense

OpenStudy (anonymous):

hartnn NO ZERO -9

hartnn (hartnn):

relax :3 now for x+5 = -4 again subtract 5 from both sides, what u get ?

OpenStudy (anonymous):

okay my bad...

OpenStudy (anna1448):

okay x+5 =-4 -5 -5 0 1< positive 1

OpenStudy (anonymous):

-4-5=??

hartnn (hartnn):

0 correct again \(\Large -4-5 = - (4+5) = ... ?\)

OpenStudy (anonymous):

it is not 5-4 dear

OpenStudy (anonymous):

again it will be a lot easier if u just show us ur calculations

OpenStudy (anna1448):

ohhh wait okay so x+5 = -4 - 5 so negative four and the difference of positive 5?

OpenStudy (anonymous):

no that mean negative four + negative 5

OpenStudy (anonymous):

WHAT UR SAYING IS THIS -4-(-5)

OpenStudy (anonymous):

which will give u 1

hartnn (hartnn):

+ negative 5 is same as difference of 5 she was correct lol

OpenStudy (anna1448):

._.

OpenStudy (anonymous):

-_-.I didnt read it

OpenStudy (anonymous):

properly.....lol

OpenStudy (anna1448):

yay i was right ? :D

hartnn (hartnn):

no ptoblem :) anyways Anna whats 4+5 ?

hartnn (hartnn):

*problem

OpenStudy (anonymous):

I am just going crazy over that problem SITH HAS GIVEN......almost solved it

OpenStudy (anna1448):

9

hartnn (hartnn):

yes so \(-4-5 = -(4+5) = -9\) makes sense ? ask doubts

OpenStudy (anna1448):

yes i think it makes sense :)

OpenStudy (anonymous):

Yesssssss........finally whats -5-6=??

OpenStudy (anna1448):

11

hartnn (hartnn):

cool:) so the answers you got are x =-1 and x=-9 \(\Large (-1,-9)\) is your final answer :)

OpenStudy (anonymous):

NO NO

hartnn (hartnn):

which is 3rd option :3

OpenStudy (anonymous):

ITS NOT 11

OpenStudy (anonymous):

U FORGOT THE SIGN

OpenStudy (anonymous):

-5-6=-(5+6)

OpenStudy (anna1448):

IM SORRY

OpenStudy (anonymous):

get it???

OpenStudy (anna1448):

-11

OpenStudy (anonymous):

OKAY NOW............I can relax........:-)

OpenStudy (anonymous):

hartnn U solved that other problem of SITH???

OpenStudy (anna1448):

im dying lol :"D

OpenStudy (anonymous):

bfore that..........what do u get ur answr??

OpenStudy (anna1448):

x =-1 and x=-9 this is it right

hartnn (hartnn):

Anna, we are done with this question, answer is 3rd option :)

hartnn (hartnn):

yes, thats it

OpenStudy (anna1448):

thank you @hartnn and @cambrige :)

hartnn (hartnn):

i had a question.... why isn't your eyes seen in your profile pic? :P

OpenStudy (anna1448):

cuz i think i am really ugly

hartnn (hartnn):

you're clearly not! i thought your were trying to show something so much that you cropped your eyes??! :P

OpenStudy (anna1448):

nah im not a slut XD

OpenStudy (anna1448):

but thnx

OpenStudy (anna1448):

and*

OpenStudy (anonymous):

yah....not interested in that...lmao

OpenStudy (anonymous):

as long as u can solve stuff.......me got no problemo

OpenStudy (anna1448):

lmao >.<

OpenStudy (anonymous):

no medals for me?? :(..

OpenStudy (anonymous):

lol........just kiddin......CAmbrige out

OpenStudy (anna1448):

ahhhhhhhh no OMG IM SO SORRY

OpenStudy (anna1448):

WAIT I FEEL SO BAD NOW :(

hartnn (hartnn):

Anna, i can give it to him, but then u won't have one :P

OpenStudy (anna1448):

i dont care :) you can give it to him :)

hartnn (hartnn):

done! :)

OpenStudy (anna1448):

yayyyy :) thnx

OpenStudy (anonymous):

guys imma kiddin

hartnn (hartnn):

cam, give her one, so we all have one!

OpenStudy (anonymous):

take the medal dude........I aint a kid yo

OpenStudy (anonymous):

that was humour............to relax meself.....

hartnn (hartnn):

yaay, we all have one now! Anna :)

OpenStudy (anonymous):

there I medalled her

OpenStudy (anna1448):

lol yayyy :D thnx man @cambrige

OpenStudy (anna1448):

and thnx @hartnn

hartnn (hartnn):

my pleasure ^_^

OpenStudy (anonymous):

np..yw.....u shud thank hartnn...he's the man!!

OpenStudy (anna1448):

i did :)